Two unbiased dice are thrown. find the probability that (i) the first die shows 6. (ii) the total of the numbers on the dice is 8. (iii) both the dice show the same number. (iv) the total of the numbers on the dice is greater than 8. (v) the total of the numbers on the dice is 13. (vi) the total of the numbers on the dice is any number from 2 and 12, both inclusive

Grade:12th pass

1 Answers

Arun
5 years ago
 
Dear student
 
(i)-
p(E) = 1/6
 
(ii)-
sum is 8 = (2,6) (5,3) (3,5)(4,4)(6,2)
 
hence p(E) = 5/36
 
similarily you can do the rest parts. If you find any difficulty please feel free to ask